This problem I and having major difficulty with. Could someone explain this in a simple way. My teacher made it very difficult and didn't clear up many of my fellow students questions.
Q: The distance of an object above the ground at time t is given by h(t) = cos(t). What is the velocity v(pi/3) of the object at t =pi/3.
Thank you!!

Answer by rothauserc(4718)

The first derivative gives us the velocity, therefore
h'(t) = -sin(t)
now t=pi/3 is in radians
h'(t) = -sin(pi/3) = 0.304810621 downwards